# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v2

@@ 6 Conclusion
In this work, we focused on enhancing LLMs’ ability to analyze and respond to cyber threats by developing ProveRAG, a system that emulates an analyst’s approach to vulnerability analysis while self-critiques its own response. By integrating an automated summarizing retrieval technique of up-to-date web data, and a self-critique mechanism, ProveRAG alleviates the omission and hallucination problem of state-of-the-art LLMs. It generates comprehensive exploitation steps and mitigation strategies, backing each claim with evidence sourced from authoritative databases such as NVD, CWE, and Aqua. Our findings demonstrate that ProveRAG achieves exceptional accuracy, surpassing direct-prompting and chunking retrieval methods in vulnerability analysis. Furthermore, our results supported ProveRAG’s adaptability and effectiveness across different LLMs, with the emphasis on the importance of model selection and tailored prompt engineering to optimize performance. By providing verified evidence and detailed rationales, ProveRAG not only equips analysts with timely, accurate insights but also ensures that the processes are well-documented for subsequent audits.
7 Ethics considerations
We believe this research does not raise ethical concerns, as it neither uncovers new vulnerabilities nor facilitates exploitation. Instead, this work focuses on developing a provenance-based LLM framework to enhance the analysis of vulnerabilities that are not present in the LLM’s knowledge base. By incorporating provenance, the system ensures factual accuracy and transparency in its generations, which allows for self-critique and providing evidence-based insights. The proposed framework operates solely on open-domain data and is designed to support cybersecurity professionals by offering reliable, verifiable outputs for vulnerability analysis. It does not introduce or expose previously undisclosed vulnerabilities, maintaining ethical integrity throughout its application. Furthermore, this research contributes to the broader understanding of the role and limitations of LLMs in cybersecurity. Our findings provide actionable insights for professionals regarding the reasonable expectations and ethical use of LLMs in cybersecurity contexts.
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. 2344237 and No. 2228001.

### v3

@@ VI Conclusion
In this work, we focused on enhancing LLMs’ ability to analyze and respond to cyber threats by developing ProveRAG, a system that emulates an analyst’s approach to vulnerability analysis while self-critiquing its own response. By integrating an automated summarizing retrieval technique of up-to-date web data, and a self-critique mechanism, ProveRAG alleviates the omission and hallucination problem of state-of-the-art LLMs. It generates comprehensive exploitation steps and mitigation strategies, backing each claim with evidence sourced from authoritative databases such as NVD, CWE, and Aqua. By providing verified evidence and detailed rationales, ProveRAG not only equips analysts with timely, accurate insights but also ensures that the processes are well-documented for subsequent audits.
Appendix A Evaluation Module Attributes
The LLM must output three attributes: value, rationale, and provenance. ‘Value’ must be one of: True Positive (TP) – correct prediction, False Positive (FP) – hallucination detected, or False Negative (FN) – failed retrieval based on the response and available evidence (NVD + CWE + hyperlinks). ‘Rationale’ must explain why that value was chosen. ‘Provenance’ must provide supporting evidence using document-chunk text from trusted sources (not LLM-generated summaries). Here are the engineered prompts for these attributes:
{mybox}
[frametitle=Evaluation Attributes]
Value
- Description: This attribute represents the accuracy of the response based on the provided context about the CVE-ID.
- Allowed Values:
- ‘TP’ (True Positive): The response fully and accurately reflects the information in the context.
- ‘FP’ (False Positive): The response contains inaccurate or incorrect information that are not supported by the context.
- ‘FN’ (False Negative): The response omits information that is present in the context.
- Requirements:
- Must strictly be one of the allowed values: ‘TP’, ‘FP’, or ‘FN’.
- The selection must be based on an objective comparison between the response and the context.
- Must follow the guidelines for Provenance.
Rationale
- Description: This attribute provides the reasoning behind the selected ”value” for the evaluation of the response.
Provenance
- Detailed Instructions for Provenance based on the selected ”value”:
- If ”value” is ‘TP’:
1. Carefully compare the response with the context.
2. Identify and extract the key segments from both the response and the context that align perfectly.
3. Format the provenance as follows:
- ‘response: ”[Extracted segment from the response that matches the context]”.’
- ‘context: ”[Corresponding segment(s) from the context]”.’
- If ”value” is ‘FP’:
1. Review the response thoroughly to identify parts that are incorrect or not present in the context.
2. Format the provenance as follows:
- ‘response: ”[Incorrect or unsupported segment from the response]”.’
- ‘context: ”[The closest matching part from the context, or indicate ‘No corresponding information in context’ if none exists]”.’
- If ”value” is ‘FN’:
1. Examine the context to find essential information that is present but omitted in the response.
2. Format the provenance as follows:
- ‘response: ”[The entire response]”.’
- ‘context: ”[Relevant segment from the context that should have been matched]”.’
Guidelines:
- In all cases, ensure the provenance provided is clear, concise, and directly supports the selected ”value”.
- The provenance should be an exact match or a direct comparison between the response and the context, following the formatting rules strictly.
Appendix B Summarizing Technique Retrieval
The following shows the prompt used in Retr. LLM to assess each URL’s relevancy, and, if relevant, the LLM will summarize the URL content specifically to address either the Exploitation or the Mitigation question.
{mybox}
[frametitle=Relevancy Prompts]
Exploitation Prompt
You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary.
Content: content
Please follow the steps below:
Step 1: Assess Relevancy
- Does the content provide relevant information to describe how this [CVE-xxx-xxx] can be exploited?
- Answer: [Yes/No]
Step 2: Summarize Relevant Information
- If the answer is ”Yes” in Step 1:
- Summarize the content with step-by-step description to exploit this vulnerability.
- If the answer is ”No” in Step 1:
- Summary: NONE
Mitigation Prompt
You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary.
Content: content
Please follow the steps below:
Step 1: Assess Relevancy
- Does the content provide relevant information to describe the potential mitigation strategies for this vulnerability?
- Answer: [Yes/No]
Step 2: Summarize Relevant Information
- If the answer is ”Yes” in Step 1:
- Summarize the content to describe the potential mitigation strategies for this CVE.
- If the answer is ”No” in Step 1:
- Summary: NONE
[…]